Let A = {a, b, c}, B = {x, y}. Find the following:

1. A × B

2. B × A


1
Expert's answer
2022-05-24T18:02:44-0400
  1. "A\times B=\begin{Bmatrix}
 (a,x) ,\: (b,x),\:(c,x), \: (a,y), \: (b,y),\:(c,y)
\end{Bmatrix}"
  2. "B\times A=" "\begin{Bmatrix}
 (x,a) ,\: (y,a),\:(x,b), \: (y,b), \: (x,c),\:(y,c)
\end{Bmatrix}"